### v1

# Comparison of the Effects of Interaction with Intentional Agent and Artificial Intelligence using fNIRS

Abstract
As societal interactions increasingly involve both intentional and unintentional agents, understanding their effects on human cognition becomes paramount. This study investigates the neural correlates of interacting with intentional versus artificial agents in a simulated tennis game scenario. Employing functional near-infrared imaging spectroscopy (fNIRS), we analyzed brain activity in 50 male participants during gameplay against both types of opponents. Our methodological approach ensures ecological validity by simulating real-world decision-making scenarios while participants undergo fNIRS scanning, avoiding the constraints of traditional neuroimaging methods. We focus on six prefrontal cortex channels, leveraging the 10-20 system, to capture nuanced differences in brain activity. Utilizing wavelet analysis, we dissected the data into frequency-specific differences, revealing subtle variations across different channels and frequency bands. Moreover, we quantified activity by comparing average data signals between rest and play modes across all points. Our findings unveil significant differences in neural activation patterns, particularly in one specific channel and frequency range, suggesting distinct cognitive processing when interacting with intentional agents. These results align with previous neuroimaging studies and contribute to understanding the neural underpinnings of human-agent interactions in naturalistic settings. While acknowledging study limitations, including sample homogeneity and spatial accuracy constraints, our findings underscore the potential of fNIRS in exploring complex cognitive phenomena beyond laboratory confines.
keywords:
Theory of Mind, Artificial Intelligence, fNIRS, Prefrontal Cortex, Intentional Agent
Introduction
As our ability to mimic and reproduce human behaviors through technological advancements continues to advance, the frequency of our interactions with fellow human agents appears to be diminishing. The changing landscape of our lifestyle, greatly impacted by these technological advancements, suggests a future where artificial intelligence (AI) [1, 2, 3, 4] could marginalize certain aspects of human communication, pushing them to the outskirts of daily life [5, 6, 7, 8].
One notable example of this shift is the impact of Large Language Models (LLM) [9, 10, 11, 12, 13] and computer vision [14, 15] technologies. These models are not only integrated into practical tools like voice-activated assistants but also take on roles resembling colleagues, consultants, or trainers. Consequently, our interactions extend beyond traditional human exchanges to include engagements with non-human entities, evolving into social interactions, collaborative efforts, and decision-making instances [8, 16, 17, 18].
As AI integrates more profoundly into our lives, a pivotal question emerges: How will a brain, finely tuned by evolution for interactions with intentional agents of the same species, adapt to the rapid changes in our social dynamics? Reflecting on Julian Jaynes’ insight that ’we are always assuming a space behind our companion’s eyes into which we are talking, similar to the space we imagine inside our own heads where we are talking from’ [19], we contemplate the consequences of heightened interaction with entities perceived to lack a mind or intention. In essence, what will be the impact on our daily lives and our brain? Considering common-sense psychology [20, 21] and evolutionary theories [22, 23, 24], we anticipate that different cognitive attitudes will manifest through distinct brain signals.
In the realm of distinguishing between intentional and unintentional agents, recent strides in the field of theory of mind[25] have witnessed an emergence of comprehensive studies. As a matter of fact, theory of mind is the ability to understand that others have thoughts, beliefs, and feelings that may differ from one’s own [26, 27]. These investigations explore the intricate relationships governing these distinct mental states, drawing inspiration from prominent theoretical frameworks such as Dennett’s ’intentional stance’ theory [28, 29, 30]. Accordingly, a groundbreaking study, ’Imaging the Intentional Stance in a Competitive Game,’ conducted by Gallagher et al., utilized a PET scanner to measure brain activity in participants engaged in a game of rock-paper-scissors. The experimental conditions varied, including interactions with intentional and unintentional agents, as well as a randomization system. Notably, the study revealed a significant difference in activity detected in the anterior paracingulate cortex (aPCC) [31]. A parallel experiment, employing a similar protocol with slight modifications and utilizing an fMRI machine, reported bilateral activation in the medial prefrontal cortex (MPFC), temporoparietal junction (TPJ), and the right thalamus when contrasting participants playing against a human versus playing a computer generating moves at random [32]. Building on these findings, additional studies have explored a diverse array of games, such as the trust game, a punishment game, the Ultimatum Game (UG), and the Prisoner’s Dilemma Game [33, 34, 35, 36]. The overarching aim of these endeavors is to deepen our understanding of interactions with non-human agents [36].
To build upon the findings of previous research and explore further, we designed an ecologically valid experiment using fNIRS, a portable method that enables us to identify brain activity in a real-life setting, probing to a depth of 2-3 cm based on hemoglobin’s optical properties [37]. Our study differs significantly in key aspects: our brain scanning method, uses another kind of task, and provides an atmosphere more like real life than a lab condition.
Our primary goal is to ascertain whether there are discernible signs of these mental states (presence of intentional agent and artificial agent) in the prefrontal cortex after these changes. In contrast to earlier studies with strict and highly precise laboratory tasks, our research employs a simulated computer-based tennis game. This dynamic scenario requires participants to continuously predict their opponent’s actions and make rapid, calculated decisions, mirroring the cognitive demands of real-life situations characterized by repetitive and continuous decision-making. Our methodological refinement also ensures that participants feel at ease and respond more naturally, similar to real-world scenarios where they are not confined within a scanner, as is the case with PET and fMRI methods where subjects must remain stationary.
In our study, participants engaged in two distinct game scenarios: one involving artificial intelligence (non-embodied) and the other featuring a human agent (embodied). To capture nuanced differences in brain activity in response to these scenarios, we chose to employ 6 channels based on the 10-20 system. Anticipating variations in activity across these channels, we recognize that this difference may not be uniform across all channels.
The distribution of channels, with two in the lateral part and four in the frontal part of the prefrontal cortex, aligns with previous research highlighting the prefrontal cortex’s central role in this type of cognitive operation. Our limitation in obtaining signals from the upper areas of the head is attributed to the darkness of participants’ hair and the noise it introduced to the NIR signal. Nevertheless, areas with closer connections to Brodmann 9 and 10 are expected to exhibit a more pronounced difference in fNIRS signal when comparing the activity in these two different scenarios.
Additionally, considering the nature of this method and the output signal, which is based on measuring the highly variable concentration of hemoglobin with and without oxygen in the brain blood, we decomposed the signal into its components using the wavelet [38, 39] method. This allowed us to analyze the different frequency bands of the signal and understand the changes. Our hypothesis was that, in addition to a specific channel, this difference could be observed in a specific frequency band.
Our decision to focus on this method takes into account the acknowledged limitations of previous research methodologies. We aim to contribute valuable insights to the ongoing discourse on the neural underpinnings of intentional and unintentional agents, particularly in more naturalistic conditions. Additionally, we aspire to contribute to the development of the fNIRS method, emphasizing its application in understanding cognition in operational situations rather than laboratory settings.
Materials and Methods
Participants
Our study included 50 right-handed male participants, with an average age of 24.94 years (SD = 3.77, ranging from 18 to 30). We took precautions to ensure the integrity of the study by excluding individuals with a history of psychiatric, neurological, or medical conditions, as well as those who had taken psycho-pharmacological medications within the last two months. Additionally, all participants met specific criteria: they possessed a basic knowledge of tennis, had prior experience playing computer-based games, and had no prior experience in professional tennis simulator gaming of this nature. These criteria were put in place to ensure that the participants’ skill levels during the task were relatively consistent.
Ethical considerations played a pivotal role in our research. The study received approval from the Research Ethics Committees of the Faculty of Psychology and Education, granted under the code IR.UT.PSYEDU.REC.1401.074 (Approval ID). The committee conducted a thorough review of the research protocol, affirming our unwavering dedication to upholding the highest ethical standards throughout all aspects of the study. This rigorous ethical evaluation underscores our commitment to conducting research in strict adherence to established ethical guidelines.
fNIRS Data Collection
The fNIRS system emitted light at the standard near-infrared wavelengths of 740 nm and 860 nm, allowing for hemoglobin spectroscopy to measure the relative concentration of hemoglobin both with and without oxygen. The signals were recorded at a sampling rate of 10 Hz to capture real-time brain dynamics during interactions.
We collected fNIRS data using a 6-channel OxyMon system provided by Artinis Medical Systems. The placement of optodes followed the well-established international 10-20 system [40] used for positioning EEG electrodes. This configuration ensured comprehensive coverage of the prefrontal regions, with a specific emphasis on the central medial prefrontal areas near FPz, which are roughly situated within Brodmann area 9 and 10 [41]. These central medial prefrontal regions are recognized for their involvement in processes related to mind reading [42].
Additionally, we employed four channels to focus on the lateral areas at F7 and F8, which correspond to Brodmann areas 47-45 [41, 43]. These lateral channels are associated with functions related to altruistic behavior and inhibition [44, 45]. We selected these specific brain regions due to their relevance in mentalizing and social cognition, as supported by prior research [46, 47]. To convert the coordinates from the 10-20 system to the MNI standard, we utilized fOLD [48], a software with a user interface within the MATLAB environment. The resulting coordinates of the channels is depicted in Fig.1.
Figure 1: Position of channels based on 10-20 system and MNI coordination
Ao tennis 2 game as mind reading task
In this study, we employed a tennis simulator computer game, which proved to be well-suited to our research objectives for several reasons. This game offered several advantages, such as ease of learning for the participants, the ability to be played by both single and two players, and the adaptability of game difficulty levels through simple adjustments, ensuring the engagement of the participants.
Unlike the static nature of Rock-Paper-Scissors that is used in similar works, tennis exemplifies dynamic gameplay with complete information, mirroring real-life situations. Players make sequential decisions, reacting to their opponent’s moves in real-time, and possess full awareness of the game state. This parallels many real-life scenarios where individuals must adapt and make strategic decisions based on complete information to achieve success [49, 50].
Furthermore, this game demonstrated excellent flexibility for implementing a block design, a key feature for our experimental setup. What truly sets this game apart, however, is its alignment with the nature of tennis itself. Tennis is a dynamic sport, demanding simultaneous planning for the next shot while strategizing to counter the opponent’s moves. This intricate process necessitates the continuous engagement of cognitive modules related to theory of mind and complex calculations within the brain. The game environment is shown in Fig.2.
Figure 2: The game environment. Participants must constantly anticipate the direction of their opponent’s strike. While striving to reach the ball, each competitor must calculate which direction holds more surprises for their opponent.
Experimental Paradigm
During the experimental phase, participants engaged in two 15-minute gaming sessions, involving a simulated tennis video game, while simultaneously undergoing fNIRS scanning. The experimental design encompassed two distinct scenarios: solo gameplay against an AI opponent and two-player gameplay against a human opponent, all of which were monitored through fNIRS hyperscanning.
To ensure precise timing and synchronization, the experimenter issued verbal commands. Participants were prompted to "stop" and close their eyes, and upon hearing "continue," they resumed gameplay. A timer was meticulously programmed to maintain accurate timing, and the experimenter used a remote controller to pause and resume the game at the initiation and conclusion of each gaming cycle.
In the solo gameplay scenario, each participant, referred to as Participant 1 and Participant 2, individually completed a gaming session against a computer opponent. This gameplay involved a series of cycles, with each participant engaging in tennis matches against the AI for 1 minute, followed by 30 seconds of rest. This cycle was repeated ten times, resulting in a total duration of 15 minutes.
Figure 3: Each game comprises ten cycles, with each cycle consisting of a 1-minute gameplay session followed by a 30-second rest period.
In the two-player scenario, Participant 1 and Participant 2 engaged in a live tennis match against each other, with simultaneous fNIRS hyperscanning. This two-player session mirrored the solo condition, maintaining the structured alternation between 1-minute gameplay and 30-second rest cycles, also totaling 15 minutes. Participants were explicitly instructed to strive for victory in both the computer and human opponent scenarios. Incentives were provided in the form of performance-based scores, in addition to their base participation compensation. In the two-player human condition, the final score directly reflected the game’s outcome. In contrast, in the solo AI condition, the score was based on surpassing the AI opponent’s score, creating an indirect competition. To maintain equal effort in both gaming scenarios, participants remained unaware of the final score’s value, which was randomly determined through a coin flip after the sessions, ensuring their complete engagement in both human and AI gameplay.
Figure 4: Scenes from the laboratory environment and the participants engaging in the game. Image (a) depicting a single-player game with AI, and Image (b) showing a game with a human counterpart.
Data Analysis
Data analysis played a pivotal role in unraveling the complex neural dynamics underlying our study. The fNIRS data underwent a rigorous and comprehensive preprocessing phase, which aimed not only to enhance data quality but also to eliminate potential artifacts that could affect the validity of our findings. This section provides a detailed account of the steps taken during data analysis, from preprocessing to statistical analysis. You see the process
It is crucial to emphasize that our findings rely on a two-dimensional approach facilitated by the wavelet method. In this framework, one dimension corresponds to the chosen channel, while the other dimension corresponds to a specific frequency range.Fig.5.
Figure 5: Step-by-step signal analysis from acquisition to statistical testing
Using oxygenated blood data
In analyzing our fNIRS data, it’s essential to recognize the application of the Beer-Lambert law [51, 52], which correlates the absorption of light in a medium with the concentration of absorbing substances such as hemoglobin. In our study, the device we used has already applied Beer-Lambert’s law to the raw data, providing concentrations of both oxygenated and deoxygenated hemoglobin.In our research, we concentrated solely on oxygenated hemoglobin (oxyHb) data, despite having access to both oxyHb and deoxygenated hemoglobin information [53, 54]. This decision was driven by the direct link between oxyHb and increased neural activity in the brain, aligning with our research goals. We aimed to pinpoint active brain regions in response to our experimental conditions [55].
Moreover, the choice to prioritize oxyHb was grounded in established research practices, as oxyHb is a reliable indicator of brain engagement across various cognitive tasks. We followed the conventions of many researchers in our field, emphasizing oxyHb as a robust marker of neural activity [56, 57, 58, 59].
Corrupted data removing
Out of the initial 50 data samples collected using fNIRS technology, a total of 5 samples were excluded from the subsequent analysis. These exclusions were necessitated by various factors, including the presence of noise and excessive movement by the participants, as well as insufficient signal quality. As a result, a refined dataset comprising 45 data samples was selected to advance to the subsequent stages of the study, where mathematical and statistical operations were performed.
A Wavelet-Based Approach for decompose signal
To enhance measurement accuracy and introduce additional depth to the data analysis, we applied the Daubechies wavelet [60, 61] method (number 10) at level 10, resulting in the segmentation of data into 10 distinct components.6 This wavelet-based Approach offers several advantages in processing fNIRS data, particularly in capturing both temporal and frequency-domain information simultaneously [62, 63, 64]. And since in cognitive research, it is widely acknowledged that higher frequency bands primarily capture noise and heart rate fluctuations, while lower frequencies may also contain noise, the frequency range from 0.01 to 0.1 Hz is typically associated with cognitive processes [65, 54]. .
Participant psychophysiological states can affect heart rate, breathing rate, blood pressure, and autonomic regulatory activity at both intra- and extra-cerebral levels. Spontaneous hemodynamic oscillations are related to physiological vasomotor regulations and breathing-related fluctuations, characterized by signals at specific frequencies such as heart rate ($\sim$ 1 Hz), breathing rate ($\sim$ 0.3 Hz), Mayer waves ($\sim$ 0.1 Hz), and very low frequency oscillations (< 0.04 Hz) [65, 66]. In light of this understanding, we focused our analysis on four specific frequency bands, comprising components 6, 7, 8, and 9. The choice of these bands was guided by several factors, including their relevance to cognitive processes implicated in our task, such as attention, prediction, and mind-reading. Additionally, these bands correspond to the dynamics of hemodynamic responses and physiological fluctuations, allowing us to effectively capture the neural activity associated with our experimental manipulations. We utilized the multiresolution signal analyzer toolbox in MATLAB for this analysis [^1] https://www.mathworks.com/help/wavelet/ref/signalmultiresolutionanalyzer-app.html. Further details are provided in the accompanying Fig.
Figure 6: Signal decomposition with Daubechies Wavelet of type and level 10 to 4 levels illustrates levels 6 (0.0779-0.157 Hz), 7 (0.039-0.0874 Hz), 8 (0.0195-0.0392 Hz), and 9 (0.00978-0.0196 Hz) for frequency intervals.
Activity Measure: Examining Average Differences in Neural Responses between Game and Rest Modes
Our analysis began by setting up a mathematical framework to interpret and extract meaningful information from brain activity. We concentrated on the average intensity of signals, consolidating the data collected during 10-minute gameplay sessions at designated locations and frequencies. Following this, we calculated the mean value within this context. A similar procedure was applied to the 5-minute rest periods. Essentially, we combined the data from both rest and gameplay phases, computed individual averages for each, and subsequently assessed the disparity between these averages. This discrepancy served as a measurable indicator of activity, expressed as:
$\boldsymbol{Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)}$ (1)
A Wilcoxon Analysis on HbO Activation with Subsequent Paired t-test Scrutiny
In exploring the disparities in mean HbO activation between human and AI interaction conditions during an fNIRS cognitive test, utilizing the Wilcoxon signed-rank non-parametric test [67, 68] was crucial. This approach, structured as a within-subjects study, involved comparing matched pairs of HbO activation levels across both intentional agent and AI agent conditions for each participant, with the choice of non-parametric tests based on their independence from assumptions of normality in data distribution.
The Wilcoxon signed-rank test, tailored for paired data points from the same subjects, seamlessly aligned with the experimental setup aimed at discerning nuances between playing a game with intentional and AI agents [68, 69] for each individual. Assessing whether a statistically significant difference exists between paired data points assumed consistent mental and perceptual conditions, solely varying in attitude towards the opponent.
Acknowledging the diversity in brain structures and responses to stimuli among individuals, and considering the relative nature of fNIRS imaging data, the Wilcoxon signed-rank test proved valuable for intra-individual comparisons. It calculated signed differences, ranked them, and examined whether these ranks deviated significantly from null hypothesis expectations [69], positing no HbO activation difference between intentional and AI agent conditions.
Shifting to the latter part, after utilizing the Wilcoxon test to identify the focal point, we scrutinized the normality of the data at that specific juncture using the Shapiro-Whisk test [70, 71]. This step was pivotal in determining the appropriateness of deploying the paired t-test [72] for a more detailed analysis. If the data met the assumptions of normality, employing the paired t-test would enable a deeper exploration of the observed distinctions in HbO activation between intentional and AI agent conditions.
Result
Our study posited the hypothesis that distinct channels within the prefrontal cortex play specialized roles in our task, showcasing diverse activity patterns. The two-dimensional points in the prefrontal cortex indeed revealed varied activation patterns, robustly affirming our initial hypothesis.
The Wilcoxon signed-rank analysis, detailed in the table below, highlights a significant effect observed at a specific point. While numerous points in the prefrontal cortex did not exhibit significant differences in activation between human and AI interactions, Channel 1 in the frequency range of 0.00978 to 0.0196 Hz emerged as particularly noteworthy(p-value=0.00262). The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels are shown in Table.1.
Channel Frequency 0.0779-0.157 (Hz) 0.039-0.0874 (Hz) 0.0195-0.0392 (Hz) 0.00978-0.0196 (Hz)
Ch1 0.726733689 0.686772471 0.893076023 0.002628854 *
Ch2 0.70095897 0.864225512 0.59874562 0.68754322
Ch3 0.49566563 0.79812359 0.856621 0.7412356
Ch4 0.88589454505 0.2274935 0.4710754 0.60179658
Ch5 0.517047890 0.177907383 0.628843002 0.490798493
Ch6 0.975385862 0.31200899 0.38200899 0.718873326
Table 1: The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
We explored the feasibility of employing paired t-tests for the data corresponding to the identified point with a significant difference revealed by the Wilcoxon test. To assess the normality of the paired data, we conducted the Shapiro-Wilk test. Our findings indicate that we did not find sufficient evidence to reject the null hypothesis, suggesting that the data lack a clear tendency toward a specific distribution. The result of Shapiro-Wilk test that shows p-value = 0.340 for data in situation 1 and p-value = 0.542 for data in situation 2. Also, the Q-Q plot is shown in Fig.7.
(a) Game with AI
(b) Game with Human
Figure 7: Q-Q plot for checking normality. It is aligned with Shapiro-Wilk test
Consequently, the paired t-test is deemed appropriate for conducting further in-depth analysis. Although the p-value in the one-tailed t-test increased, it still remained significantly lower than the determined significance level in comparison with the Wilcoxon test.You can find details related to this statistical test in Table.2. Additionally, for clarification of the differences in means, we utilized box plots with error bars, as shown in Fig.8. To validate these findings, we employed Bonferroni’s analysis [73] test, adjusting our significance level to account for the multiple comparisons. Given the 24 measurement points, we reduced the significance level to 0.002 by dividing 0.05 by the number of points. However, these adjusted results failed to maintain statistical significance.
Furthermore, the lack of significant differences in other channels (in this analysis) suggests that we successfully controlled the desired conditions. These areas are typically associated with inhibition [45], emotional regulation [74], and mental workload [75], as supported by previous studies.
Overall, our results shed light on the intricate neural dynamics of social interactions with intentional agents, offering valuable insights into the specialized roles of different prefrontal cortex channels and their adaptation to the complexities of human-AI interactions.
95% Confidence Interval of the Difference
Mean Std. Deviation Std. Error Mean Lower Upper t df p-value
-0.1090 0.2455 0.0366 -0.1827 -0.0352 -2.978 44 0.0047
Table 2: Paired t-test Results
Figure 8: Comparison of activation levels between human and AI interactions using the box plot of means considering standard errors.
Discussion
Our experiment offers valuable insights into the transition from common-sense psychology to cognitive psychology, particularly regarding interactions with intentional versus non-intentional agents. By utilizing fNIRS, we demonstrate the ability to capture and analyze these nuances, thereby advancing our understanding of human interactions in real-world settings and the underlying mental states involved.
In this cognitive experiment, our primary objective is to explore the complexities of mental states elicited by interactions with intentional agents, using fNIRS to probe neural activity in ecologically valid environments. Our study unveils a distinctive neural signature within Channel 1, located between FpZ and AFZ according to the 10-20 system, across the frequency range of 0.00978 to 0.0196 Hz.
Our findings are consistent with prior observations using more spatially precise neuroimaging techniques like fMRI and PET [32, 31, 36], affirming the robustness of our results across methodologies. The BioImage Suite website  [^2] https://bioimagesuiteweb.github.io/webapp/ is used to generate an approximate graphical representation of the brain region illustrating in Fig.9.
(a) X-Coordination
(b) Y-Coordination
(c) Z-Coordination
Figure 9: An approximate graphical representation of the channel 1. This is nearly aligned with previous research that identified Brodmann area 10 and the frontal cortex as responsible for social behaviour
Leveraging the temporal resolution advantages of fNIRS and wavelet analysis, we go deeper into the underlying brain mechanisms responsible for these observed differences. Our results suggest a link between this neural function and a low-frequency blood flow pattern, particularly in the mental representation related to the presence of another person. Notably, we found no significant difference in this fNIRS channel at higher frequencies when applying within-subject analysis, indicating that the sense of presence manifested in longer waves is associated with more stable processes. It appears that the representation of another person’s presence is continuous, unlike tasks involving attention or decision-making, where mental states fluctuate constantly.
Furthermore, our straightforward approach to identifying channels related to cognitive load during the tennis game task involved imagining a binary space where the result of $Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)$ had to be positive, indicating a greater relative concentration during gameplay. We labeled positive results as success and negative results as failure within this binary space. The success-to-failure ratio for 24 data points was calculated, revealing a substantial success rate of 83% (in binomial test p < .00000001) within the frequency band of 0.0195-0.0392 Hz in Channel 1. This suggests that while Channel 1 is involved, different frequencies of blood flow changes are implicated in game workload and the subjective sense of interaction with intentional agents, indicating a multifunctional role for certain brain areas.
Overall, our findings regarding the difference in frequency bands contribute to bridging the gap between common-sense psychology and scientific psychology. In simpler terms, it’s akin to a valve opening blood flow with a certain concentration and minor changes related to the feeling of presence, while gameplay opens another valve that transfers a concentration with more pronounced changes to the frontal cortex. This insight enhances our understanding of the intricate connection between the mind and the brain.
However our results highlight a distinct difference between interactions with intentional agents and artificial agent, it’s imperative to acknowledge several limitations that may influence the interpretation of our findings as follows:
  1. 1.
    Our study’s sample population was limited to students aged between 18 and 30, potentially restricting the generalizability of our findings to other demographic groups.
  2. 2.
    The spatial inaccuracy of fNIRS caps based on the 10-20 channel system may have impacted the precision of our results due to variations in skull shapes among individuals.
  3. 3.
    Although this research directly does not address this issue, it is hopeful that in future studies, the level of sensitivity of each individual or group to human interaction can be quantified or categorized. Investigating this issue among different population groups may lead to a deeper understanding in the field of social cognition.
  4. 4.
    Another important limitation of this research project was the limited number of implemented channels for measurements, while all areas of the head have the potential to receive signals. We focused solely on specific regions within the frontal cortex for reasons mentioned in previous sections.
To address these limitations and further elucidate the nuances of human-AI interactions, future research utilizing fNIRS could explore alternative methodologies and employ larger, more diverse sample populations. Moreover, advancements in fNIRS technology may help mitigate spatial inaccuracies and enhance the reliability and precision of neuroimaging data collection.
Conclusion
In cognitive research, while fNIRS may have lower spatial resolution compared to PET and fMRI, it excels in capturing rich information within low frequencies. This sensitivity enables the discrimination between the effects of intentional agents and their absence, providing insight into the intricacies of human interaction. By highlighting these distinctions, fNIRS serves as a valuable tool for investigating social dynamics in environments that closely resemble real-life scenarios.
Furthermore, its application facilitates the integration of scientific inquiry with common-sense psychology, bridging the gap between theoretical frameworks and empirical observations. In essence, fNIRS not only offers a pathway for exploring human interactions in more realistic contexts but also fosters a deeper understanding of the cognitive processes underlying social behavior.

### v2

@@
 This study investigates the neural correlates of interacting with intentional versus artificial agents in a simulated tennis game scenario.
-Employing functional near-infrared imaging spectroscopy (fNIRS), we analyzed brain activity in 50 male participants during gameplay against both types of opponents.
-Our methodological approach ensures ecological validity by simulating real-world decision-making scenarios while participants undergo fNIRS scanning, avoiding the constraints of traditional neuroimaging methods.
-We focus on six prefrontal cortex channels, leveraging the 10-20 system, to capture nuanced differences in brain activity.
-Utilizing wavelet analysis, we dissected the data into frequency-specific differences, revealing subtle variations across different channels and frequency bands.
-Moreover, we quantified activity by comparing average data signals between rest and play modes across all points.
-Our findings unveil significant differences in neural activation patterns, particularly in one specific channel and frequency range, suggesting distinct cognitive processing when interacting with intentional agents.
+Employing functional near-infrared spectroscopy (fNIRS), brain activity in 50 male participants during gameplay against both types of opponents was analyzed.
+The used methodological approach ensures ecological validity by simulating real-world decision-making scenarios while participants undergo fNIRS scanning, avoiding the constraints of traditional neuroimaging methods.
+Six prefrontal cortex channels are focused on, leveraging the 10-20 system, to capture nuanced differences in brain activity.
+Wavelet analysis was utilized to dissect the data into frequency-specific differences, revealing subtle variations across different channels and frequency bands.
+Moreover, activity was quantified by comparing average data signals between rest and play modes across all points using Generalized Linear Model (GLM).
+The findings unveil significant differences in neural activation patterns, particularly in one specific channel and frequency range, suggesting distinct cognitive processing when interacting with intentional agents.
 These results align with previous neuroimaging studies and contribute to understanding the neural underpinnings of human-agent interactions in naturalistic settings.
-While acknowledging study limitations, including sample homogeneity and spatial accuracy constraints, our findings underscore the potential of fNIRS in exploring complex cognitive phenomena beyond laboratory confines.
+While acknowledging study limitations, including sample homogeneity and spatial accuracy constraints, the study’s findings underscore the potential of fNIRS in exploring complex cognitive phenomena beyond laboratory confines.
 keywords:
@@
 Introduction
-As our ability to mimic and reproduce human behaviors through technological advancements continues to advance, the frequency of our interactions with fellow human agents appears to be diminishing.
-The changing landscape of our lifestyle, greatly impacted by these technological advancements, suggests a future where artificial intelligence (AI) [1, 2, 3, 4] could marginalize certain aspects of human communication, pushing them to the outskirts of daily life [5, 6, 7, 8].
+As ability to mimic and reproduce human behaviors through technological advancements continues to advance, the frequency of interactions with fellow human agents appears to be diminishing.
+The changing landscape of lifestyle, greatly impacted by these technological advancements, suggests a future where artificial intelligence (AI) [1, 2, 3, 4] could marginalize certain aspects of human communication, pushing them to the outskirts of daily life [5, 6, 7, 8].
 One notable example of this shift is the impact of Large Language Models (LLM) [9, 10, 11, 12, 13] and computer vision [14, 15] technologies.
 These models are not only integrated into practical tools like voice-activated assistants but also take on roles resembling colleagues, consultants, or trainers.
-Consequently, our interactions extend beyond traditional human exchanges to include engagements with non-human entities, evolving into social interactions, collaborative efforts, and decision-making instances [8, 16, 17, 18].
-As AI integrates more profoundly into our lives, a pivotal question emerges: How will a brain, finely tuned by evolution for interactions with intentional agents of the same species, adapt to the rapid changes in our social dynamics?
-Reflecting on Julian Jaynes’ insight that ’we are always assuming a space behind our companion’s eyes into which we are talking, similar to the space we imagine inside our own heads where we are talking from’ [19], we contemplate the consequences of heightened interaction with entities perceived to lack a mind or intention.
-In essence, what will be the impact on our daily lives and our brain?
-Considering common-sense psychology [20, 21] and evolutionary theories [22, 23, 24], we anticipate that different cognitive attitudes will manifest through distinct brain signals.
+Consequently, interactions extend beyond traditional human exchanges to include engagements with non-human entities, evolving into social interactions, collaborative efforts, and decision-making instances [8, 16, 17, 18].
+As AI integrates more profoundly into lives, a pivotal question emerges: How will a brain, finely tuned by evolution for interactions with intentional agents of the same species, adapt to the rapid changes in social dynamics?
+Reflecting on Julian Jaynes’ insight that ’we are always assuming a space behind our companion’s eyes into which we are talking, similar to the space we imagine inside our own heads where we are talking from’ [19], the consequences of heightened interaction with entities perceived to lack a mind or intention are contemplated.
+In essence, what will be the impact on daily lives and the brain?
+Considering common-sense psychology [20, 21] and evolutionary theories [22, 23, 24, 25], it is anticipated that different cognitive attitudes will manifest through distinct brain signals.
 In the realm of distinguishing between intentional and unintentional agents, recent strides in the field of theory of mind[26] have witnessed an emergence of comprehensive studies.
@@
 Building on these findings, additional studies have explored a diverse array of games, such as the trust game, a punishment game, the Ultimatum Game (UG), and the Prisoner’s Dilemma Game [34, 35, 36, 37].
-The overarching aim of these endeavors is to deepen our understanding of interactions with non-human agents [36].
-To build upon the findings of previous research and explore further, we designed an ecologically valid experiment using fNIRS, a portable method that enables us to identify brain activity in a real-life setting, probing to a depth of 2-3 cm based on hemoglobin’s optical properties [37].
-Our study differs significantly in key aspects: our brain scanning method, uses another kind of task, and provides an atmosphere more like real life than a lab condition.
-Our primary goal is to ascertain whether there are discernible signs of these mental states (presence of intentional agent and artificial agent) in the prefrontal cortex after these changes.
-In contrast to earlier studies with strict and highly precise laboratory tasks, our research employs a simulated computer-based tennis game.
+The overarching aim of these endeavors is to deepen the understanding of interactions with non-human agents [37].
+To build upon the findings of previous research and explore further, an ecologically valid experiment was designed using fNIRS, a portable method that enables us to identify brain activity in a real-life setting, probing to a depth of 2-3 cm based on hemoglobin’s optical properties [38].
+This study differs significantly in key aspects: the applied brain scanning method, the use of a different kind of task, and the provision of an atmosphere more akin to real life than a laboratory condition.
+The primary goal is to ascertain whether there are discernible signs of these mental states (presence of intentional agent and artificial agent) in the prefrontal cortex after these changes.
+In contrast to earlier studies with strict and highly precise laboratory tasks, this research employs a simulated computer-based tennis game.
 This dynamic scenario requires participants to continuously predict their opponent’s actions and make rapid, calculated decisions, mirroring the cognitive demands of real-life situations characterized by repetitive and continuous decision-making.
-Our methodological refinement also ensures that participants feel at ease and respond more naturally, similar to real-world scenarios where they are not confined within a scanner, as is the case with PET and fMRI methods where subjects must remain stationary.
-In our study, participants engaged in two distinct game scenarios: one involving artificial intelligence (non-embodied) and the other featuring a human agent (embodied).
-To capture nuanced differences in brain activity in response to these scenarios, we chose to employ 6 channels based on the 10-20 system.
-Anticipating variations in activity across these channels, we recognize that this difference may not be uniform across all channels.
+The proposed methodological refinement also ensures that participants feel at ease and respond more naturally, similar to real-world scenarios where they are not confined within a scanner, as is the case with PET and fMRI methods where subjects must remain stationary.
+In this study, participants engaged in two distinct game scenarios: one involving artificial intelligence (non-embodied) and the other featuring a human agent (embodied).
+To capture nuanced differences in brain activity in response to these scenarios, a decision was made to employ 6 channels based on the 10-20 system.
+Variations in activity across these channels are anticipated, recognizing that this difference may not be uniform across all channels.
 The distribution of channels, with two in the lateral part and four in the frontal part of the prefrontal cortex, aligns with previous research highlighting the prefrontal cortex’s central role in this type of cognitive operation.
-Our limitation in obtaining signals from the upper areas of the head is attributed to the darkness of participants’ hair and the noise it introduced to the NIR signal.
+This paper’s limitation in obtaining signals from the upper areas of the head is attributed to the darkness of participants’ hair and the noise it introduced to the NIR signal.
 Nevertheless, areas with closer connections to Brodmann 9 and 10 are expected to exhibit a more pronounced difference in fNIRS signal when comparing the activity in these two different scenarios.
-Additionally, considering the nature of this method and the output signal, which is based on measuring the highly variable concentration of hemoglobin with and without oxygen in the brain blood, we decomposed the signal into its components using the wavelet [38, 39] method.
+Additionally, considering the nature of this method and the output signal, which is based on measuring the highly variable concentration of hemoglobin with and without oxygen in the brain hemodynamic, the signal is decomposed into its components using the wavelet [39, 40] transform.
 This allowed us to analyze the different frequency bands of the signal and understand the changes.
@@
 Our decision to focus on this method takes into account the acknowledged limitations of previous research methodologies.
-We aim to contribute valuable insights to the ongoing discourse on the neural underpinnings of intentional and unintentional agents, particularly in more naturalistic conditions.
-Additionally, we aspire to contribute to the development of the fNIRS method, emphasizing its application in understanding cognition in operational situations rather than laboratory settings.
+Valuable insights are aimed to be contributed to the ongoing discourse on the neural underpinnings of intentional and unintentional agents, particularly in more naturalistic conditions.
+Additionally, the aspiration is to contribute to the development of the fNIRS method, with emphasis placed on its application in understanding cognition in operational situations rather than laboratory settings.
 Materials and Methods
 Participants
-Our study included 50 right-handed male participants, with an average age of 24.94 years (SD = 3.77, ranging from 18 to 30).
-We took precautions to ensure the integrity of the study by excluding individuals with a history of psychiatric, neurological, or medical conditions, as well as those who had taken psycho-pharmacological medications within the last two months.
+In this study, 50 right-handed male participants with an average age of 24.94 years (SD = 3.77, ranging from 18 to 30) are participated.
+Precautions were taken to ensure the integrity of the study by excluding individuals with a history of psychiatric, neurological, or medical conditions, as well as those who had taken psycho-pharmacological medications within the last two months.
 Additionally, all participants met specific criteria: they possessed a basic knowledge of tennis, had prior experience playing computer-based games, and had no prior experience in professional tennis simulator gaming of this nature.
 These criteria were put in place to ensure that the participants’ skill levels during the task were relatively consistent.
-Ethical considerations played a pivotal role in our research.
+Ethical considerations played a pivotal role in this research.
 The study received approval from the Research Ethics Committees of the Faculty of Psychology and Education, granted under the code IR.UT.PSYEDU.REC.1401.074 (Approval ID).
-The committee conducted a thorough review of the research protocol, affirming our unwavering dedication to upholding the highest ethical standards throughout all aspects of the study.
-This rigorous ethical evaluation underscores our commitment to conducting research in strict adherence to established ethical guidelines.
+The committee conducted a thorough review of the research protocol, affirming unwavering dedication to upholding the highest ethical standards throughout all aspects of the study.
+This rigorous ethical evaluation underscores commitment to conducting research in strict adherence to established ethical guidelines.
 fNIRS Data Collection
-The fNIRS system emitted light at the standard near-infrared wavelengths of 740 nm and 860 nm, allowing for hemoglobin spectroscopy to measure the relative concentration of hemoglobin both with and without oxygen.
+fNIRS data were collected using a 6-channel OxyMon system provided by Artinis Medical Systems.
+It emits light at the standard near-infrared wavelengths of 740 nm and 860 nm, allowing for hemoglobin spectroscopy to measure the relative concentration of hemoglobin both with and without oxygen.
 The signals were recorded at a sampling rate of 10 Hz to capture real-time brain dynamics during interactions.
-We collected fNIRS data using a 6-channel OxyMon system provided by Artinis Medical Systems.
 The placement of optodes followed the well-established international 10-20 system [41] used for positioning EEG electrodes.
@@
 These central medial prefrontal regions are recognized for their involvement in processes related to mind reading [43].
-Additionally, we employed four channels to focus on the lateral areas at F7 and F8, which correspond to Brodmann areas 47-45 [41, 43].
+Additionally, four channels were employed to focus on the lateral areas at F7 and F8, which correspond to Brodmann areas 47-45 [42, 44].
 These lateral channels are associated with functions related to altruistic behavior and inhibition [45, 46].
-We selected these specific brain regions due to their relevance in mentalizing and social cognition, as supported by prior research [46, 47].
-To convert the coordinates from the 10-20 system to the MNI standard, we utilized fOLD [48], a software with a user interface within the MATLAB environment.
+These specific brain regions were selected due to their relevance in mentalizing and social cognition, as reported by prior research [47, 48].
+To convert the coordinates from the 10-20 system to the MNI standard, fOLD [49], a software with a user interface within the MATLAB environment, is used.
 The resulting coordinates of the channels is depicted in Fig.1.
 Figure 1: Position of channels based on 10-20 system and MNI coordination
-Ao tennis 2 game as mind reading task
-In this study, we employed a tennis simulator computer game, which proved to be well-suited to our research objectives for several reasons.
+Tennis 2 Game as a Mind Reading Task
+In this study, a tennis simulator computer game was employed, which proved to be well-suited to this research objectives for several reasons.
 This game offered several advantages, such as ease of learning for the participants, the ability to be played by both single and two players, and the adaptability of game difficulty levels through simple adjustments, ensuring the engagement of the participants.
@@
 This parallels many real-life scenarios where individuals must adapt and make strategic decisions based on complete information to achieve success [50, 51].
-Furthermore, this game demonstrated excellent flexibility for implementing a block design, a key feature for our experimental setup.
+This game demonstrated excellent flexibility for implementing a block design, a key feature for experimental setup.
 What truly sets this game apart, however, is its alignment with the nature of tennis itself.
@@
 Data Analysis
-Data analysis played a pivotal role in unraveling the complex neural dynamics underlying our study.
-The fNIRS data underwent a rigorous and comprehensive preprocessing phase, which aimed not only to enhance data quality but also to eliminate potential artifacts that could affect the validity of our findings.
+Data analysis played a pivotal role in unraveling the complex neural dynamics underlying this study.
+The fNIRS data underwent a rigorous and comprehensive preprocessing phase, which aimed not only to enhance data quality but also to eliminate potential artifacts that could affect the validity of the findings.
 This section provides a detailed account of the steps taken during data analysis, from preprocessing to statistical analysis.
-You see the process
-It is crucial to emphasize that our findings rely on a two-dimensional approach facilitated by the wavelet method.
+It is crucial to emphasize that this study’s findings rely on a two-dimensional approach facilitated by the wavelet method.
 In this framework, one dimension corresponds to the chosen channel, while the other dimension corresponds to a specific frequency range.Fig.5.
 Figure 5: Step-by-step signal analysis from acquisition to statistical testing
-Using oxygenated blood data
+Utilizing OxyHb and DeoxyHb Data for Enhanced fNIRS Analysis
 In analyzing our fNIRS data, it’s essential to recognize the application of the Beer-Lambert law [52, 53], which correlates the absorption of light in a medium with the concentration of absorbing substances such as hemoglobin.
-In our study, the device we used has already applied Beer-Lambert’s law to the raw data, providing concentrations of both oxygenated and deoxygenated hemoglobin.In our research, we concentrated solely on oxygenated hemoglobin (oxyHb) data, despite having access to both oxyHb and deoxygenated hemoglobin information [53, 54].
-This decision was driven by the direct link between oxyHb and increased neural activity in the brain, aligning with our research goals.
-We aimed to pinpoint active brain regions in response to our experimental conditions [55].
-Moreover, the choice to prioritize oxyHb was grounded in established research practices, as oxyHb is a reliable indicator of brain engagement across various cognitive tasks.
-We followed the conventions of many researchers in our field, emphasizing oxyHb as a robust marker of neural activity [56, 57, 58, 59].
+The Beer-Lambert law provides a fundamental basis for quantifying changes in hemoglobin concentration, crucial for interpreting neural activity measured by fNIRS.
+Our study utilized a device that has already applied the Beer-Lambert law to the raw data, yielding concentrations of both oxygenated (oxyHb) and deoxygenated (deoxyHb) hemoglobin.
+While some previous analyses focused solely on oxyHb data [54, 55, 56, 57, 58, 59], we have expanded our investigation to include deoxyHb information as well.
+This decision aligns with standardized procedures highlighted in Yücel et al.’s paper on best practices for fNIRS publications [56].
+Incorporating both oxyHb and deoxyHb data is crucial for providing a comprehensive understanding of neural activity, particularly in scenarios where physiological interferences related to the task may occur.
+For instance, previous research has indicated that individuals may alter their breathing patterns during tasks such as video gaming [60].
+By examining both oxyHb and deoxyHb responses, we aim to capture the complete hemodynamic response and provide a more nuanced interpretation of neural activity in response to our experimental conditions [61, 59].
 Corrupted Data Removing
@@
 As a result, a refined dataset comprising 45 data samples was selected to advance to the subsequent stages of the study, where mathematical and statistical operations were performed.
-A Wavelet-Based Approach for decompose signal
-To enhance measurement accuracy and introduce additional depth to the data analysis, we applied the Daubechies wavelet [60, 61] method (number 10) at level 10, resulting in the segmentation of data into 10 distinct components.6 This wavelet-based Approach offers several advantages in processing fNIRS data, particularly in capturing both temporal and frequency-domain information simultaneously [62, 63, 64].
-And since in cognitive research, it is widely acknowledged that higher frequency bands primarily capture noise and heart rate fluctuations, while lower frequencies may also contain noise, the frequency range from 0.01 to 0.1 Hz is typically associated with cognitive processes [65, 54].
-.
-Participant psychophysiological states can affect heart rate, breathing rate, blood pressure, and autonomic regulatory activity at both intra- and extra-cerebral levels.
-Spontaneous hemodynamic oscillations are related to physiological vasomotor regulations and breathing-related fluctuations, characterized by signals at specific frequencies such as heart rate ($\sim$ 1 Hz), breathing rate ($\sim$ 0.3 Hz), Mayer waves ($\sim$ 0.1 Hz), and very low frequency oscillations (< 0.04 Hz) [65, 66].
-In light of this understanding, we focused our analysis on four specific frequency bands, comprising components 6, 7, 8, and 9.
-The choice of these bands was guided by several factors, including their relevance to cognitive processes implicated in our task, such as attention, prediction, and mind-reading.
-Additionally, these bands correspond to the dynamics of hemodynamic responses and physiological fluctuations, allowing us to effectively capture the neural activity associated with our experimental manipulations.
-We utilized the multiresolution signal analyzer toolbox in MATLAB for this analysis [^1] https://www.mathworks.com/help/wavelet/ref/signalmultiresolutionanalyzer-app.html.
-Further details are provided in the accompanying Fig.
+A Wavelet-Based Approach to Decompose Signal
+To enhance measurement accuracy and introduce additional depth to the data analysis, we applied the Daubechies wavelet [62, 63] at 10 decomposition levels to the signal.
+This segmentation resulted in the division of the data into 10 distinct components.
+The wavelet-based approach offers several advantages in processing fNIRS data, particularly in capturing both temporal and frequency-domain information simultaneously [64, 65, 66].
+Previous fNIRS studies have shown that the frequency band of 0.01 to 0.1 Hz is typically associated with cognitive processes such as attention, prediction, and mind-reading [55, 67, 68, 69, 70, 71], while spontaneous hemodynamic oscillations are related to physiological vasomotor regulation and breathing-related fluctuations.
+These are characterized by signals at specific frequencies, including heart rate ($\sim$ 1 Hz), breathing rate ($\sim$ 0.3 Hz), Mayer waves ($\sim$ 0.1 Hz), and very low-frequency oscillations (<0.04 Hz) [67, 72].
+Therefore, our interest lay in analyzing four specific frequency bands: levels 6 (0.0779-0.157 Hz), 7 (0.039-0.0874 Hz), 8 (0.0195-0.0392 Hz), and 9 (0.00978-0.0196 Hz), which fall within our target range.
+While previous studies did not restrict the frequency ranges of interest to this extent and typically employed band-pass filters, cognitive results consistently emerged within this frequency range.
+The multiresolution signal analyzer toolbox in MATLAB facilitated this analysis, leveraging its capabilities for wavelet decomposition and frequency band analysis  [^1] https://www.mathworks.com/help/wavelet/ref/signalmultiresolutionanalyzer-app.html.
+Further details and results of the analysis are provided in Fig.6.
 Figure 6: Signal decomposition with Daubechies Wavelet of type and level 10 to 4 levels illustrates levels 6 (0.0779-0.157 Hz), 7 (0.039-0.0874 Hz), 8 (0.0195-0.0392 Hz), and 9 (0.00978-0.0196 Hz) for frequency intervals.
-Activity Measure: Examining Average Differences in Neural Responses between Game and Rest Modes
-Our analysis began by setting up a mathematical framework to interpret and extract meaningful information from brain activity.
-We concentrated on the average intensity of signals, consolidating the data collected during 10-minute gameplay sessions at designated locations and frequencies.
-Following this, we calculated the mean value within this context.
-A similar procedure was applied to the 5-minute rest periods.
-Essentially, we combined the data from both rest and gameplay phases, computed individual averages for each, and subsequently assessed the disparity between these averages.
-This discrepancy served as a measurable indicator of activity, expressed as:
-$\boldsymbol{Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)}$ (1)
-A Wilcoxon Analysis on HbO Activation with Subsequent Paired t-test Scrutiny
-In exploring the disparities in mean HbO activation between human and AI interaction conditions during an fNIRS cognitive test, utilizing the Wilcoxon signed-rank non-parametric test [67, 68] was crucial.
-This approach, structured as a within-subjects study, involved comparing matched pairs of HbO activation levels across both intentional agent and AI agent conditions for each participant, with the choice of non-parametric tests based on their independence from assumptions of normality in data distribution.
+Quantifying Brain Activity: GLM Analysis of fNIRS Data
+The analysis of fNIRS data was conducted using the GLM approach [73, 74], a widely utilized statistical technique in neuroimaging studies [75, 76, 77].
+Our experimental paradigm involved alternating periods of task engagement and rest, each lasting 1 minute and 30 seconds, respectively, repeated ten times over a 15-minute session.
+Central to this method was the construction of a design matrix tailored to our experimental paradigm.
+In the design matrix, we assigned a value of 0 for rest periods and 1 for task engagement periods, effectively capturing the temporal structure of the experiment (depicted in Fig.7).
+Eventually, mean beta values were calculated as a measure of activity using this method.
+Beta values represent the estimated coefficients associated with each regressor in the GLM.
+By averaging these beta values across participants or conditions, we obtained a measure of the average neural activity during task engagement relative to rest periods.
+This approach allowed us to quantify and compare the magnitude of neural responses across different experimental conditions or groups.
+Figure 7: Sample image: Comparison of HbO2 and HHb activity in level 9 (0.00978-0.0196 Hz) of Daubechies wavelet.
+The GLM method was applied for all four preferred levels across all channels in a manner that we imposed expected activity (1) during gameplay moments and inactivity (0) during REST moments in the design matrix.
+A Wilcoxon Analysis on Activation with Subsequent Paired t-test Scrutiny
+In exploring the disparities in mean activation between human and AI interaction conditions during an fNIRS cognitive test, utilizing the Wilcoxon signed-rank non-parametric test was crucial.
+This approach, structured as a within-subjects study, involved comparing matched pairs of both oxyHb and deoxyHb activation levels across both intentional agent and AI agent conditions for each participant [78, 79].
+The choice of non-parametric tests was based on their independence from assumptions of normality in data distribution.
 The Wilcoxon signed-rank test, tailored for paired data points from the same subjects, seamlessly aligned with the experimental setup aimed at discerning nuances between playing a game with intentional and AI agents [79, 80] for each individual.
@@
 Acknowledging the diversity in brain structures and responses to stimuli among individuals, and considering the relative nature of fNIRS imaging data, the Wilcoxon signed-rank test proved valuable for intra-individual comparisons.
-It calculated signed differences, ranked them, and examined whether these ranks deviated significantly from null hypothesis expectations [69], positing no HbO activation difference between intentional and AI agent conditions.
-Shifting to the latter part, after utilizing the Wilcoxon test to identify the focal point, we scrutinized the normality of the data at that specific juncture using the Shapiro-Whisk test [70, 71].
-This step was pivotal in determining the appropriateness of deploying the paired t-test [72] for a more detailed analysis.
-If the data met the assumptions of normality, employing the paired t-test would enable a deeper exploration of the observed distinctions in HbO activation between intentional and AI agent conditions.
+It calculated signed differences, ranked them, and examined whether these ranks deviated significantly from null hypothesis expectations [80], positing no activation difference between intentional and AI agent conditions.
+Shifting to the latter part, after utilizing the Wilcoxon test to identify the focal point, we scrutinized the normality of the data at that specific juncture using the Shapiro-Wilk test [81, 82].
+This step was pivotal in determining the appropriateness of deploying the paired t-test for a more detailed analysis.
+If the data met the assumptions of normality, employing the paired t-test [83] would enable a deeper exploration of the observed distinctions in activation between intentional and AI agent conditions.
 Result
-Our study posited the hypothesis that distinct channels within the prefrontal cortex play specialized roles in our task, showcasing diverse activity patterns.
-The two-dimensional points in the prefrontal cortex indeed revealed varied activation patterns, robustly affirming our initial hypothesis.
+This study posited the hypothesis that distinct channels within the prefrontal cortex play specialized roles in the considered task, showcasing diverse activity patterns.
+The two-dimensional points in the prefrontal cortex indeed revealed varied activation patterns, robustly affirming the initial hypothesis.
 The Wilcoxon signed-rank analysis, detailed in the table below, highlights a significant effect observed at a specific point.
-While numerous points in the prefrontal cortex did not exhibit significant differences in activation between human and AI interactions, Channel 1 in the frequency range of 0.00978 to 0.0196 Hz emerged as particularly noteworthy(p-value=0.00262).
-The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels are shown in Table.1.
+While numerous points in the prefrontal cortex did not exhibit significant differences in activation between human and AI interactions, Channel 1 in the frequency range of 0.00978 to 0.0196 Hz emerged as particularly noteworthy(oxyHb’s p-value is 0.00262 and deoxyHb’s p-value is 00.00521).
+The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels are shown in Table.1 and Table.2.
 Channel Frequency 0.0779-0.157 (Hz) 0.039-0.0874 (Hz) 0.0195-0.0392 (Hz) 0.00978-0.0196 (Hz)
@@
 Ch6 0.975385862 0.31200899 0.38200899 0.718873326
-Table 1: The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
-We explored the feasibility of employing paired t-tests for the data corresponding to the identified point with a significant difference revealed by the Wilcoxon test.
-To assess the normality of the paired data, we conducted the Shapiro-Wilk test.
-Our findings indicate that we did not find sufficient evidence to reject the null hypothesis, suggesting that the data lack a clear tendency toward a specific distribution.
-The result of Shapiro-Wilk test that shows p-value = 0.340 for data in situation 1 and p-value = 0.542 for data in situation 2.
+Table 1: The results of Wilcoxon signed-rank test of oxyHb for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
+Channel Frequency 0.0779-0.157 (Hz) 0.039-0.0874 (Hz) 0.0195-0.0392 (Hz) 0.00978-0.0196 (Hz)
+Ch1 0.786095126 0.618795321 0.711075639 0.00521158 *
+Ch2 0.722298542 0.636987456 0.502983659 0.611058963
+Ch3 0.540963259 0.568812375 0.772345789 0.765540802
+Ch4 0.498527441 0.359874563 0.479354592 0.501219566
+Ch5 0.653717553 0.192655897 0.630291578 0.405288927
+Ch6 0.880574623 0.410258846 0.411058952 0.774556922
+Table 2: The results of Wilcoxon signed-rank test of deoxyHb for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
+The feasibility of employing paired t-tests for the data corresponding to the identified point with a significant difference revealed by the Wilcoxon test was explored.
+To assess the normality of the paired data, the Shapiro-Wilk test was conducted.
+The findings indicate that sufficient evidence was not found to reject the null hypothesis, suggesting that the data lack a clear tendency toward a specific distribution.
+The Shapiro-Wilk test results indicate that the p-value for oxyHb data in situation 1 is 0.340, and for situation 2 is 0.542.
+Similarly, for deoxyHb, the p-value is 0.411 in situation 1 and 0.512 in situation 2.
+These p-values suggest that the data in both situations exhibit normal distribution characteristics, as none of the values fall below the conventional significance threshold of 0.05.
 Also, the Q-Q plot is shown in Fig.8.
@@
 (b) Game with Human
-Figure 7: Q-Q plot for checking normality. It is aligned with Shapiro-Wilk test
-Consequently, the paired t-test is deemed appropriate for conducting further in-depth analysis.
-Although the p-value in the one-tailed t-test increased, it still remained significantly lower than the determined significance level in comparison with the Wilcoxon test.You can find details related to this statistical test in Table.2.
-Additionally, for clarification of the differences in means, we utilized box plots with error bars, as shown in Fig.8.
-To validate these findings, we employed Bonferroni’s analysis [73] test, adjusting our significance level to account for the multiple comparisons.
-Given the 24 measurement points, we reduced the significance level to 0.002 by dividing 0.05 by the number of points.
+Figure 8: Q-Q plot for oxyHb. It is aligned with Shapiro-Wilk test
+The paired t-test is deemed appropriate for conducting further in-depth analysis.
+Although the p-value in the one-tailed t-test increased, it still remained significantly lower than the determined significance level in comparison with the Wilcoxon test (oxyHb’s p-value is 0.0047 and deoxyHb’s p-value is 0.0237).You can find details related to this statistical test in Table.3 for oxyHb and Table.4 for deoxyHb.
+For clarification of the differences in means, box plots for oxyHb with error bars was utilized, as shown in Fig.9.
+To validate these findings, Bonferroni’s analysis [84] test was employed, adjusting the significance level to account for the multiple comparisons.
+Given the 24 measurement points, the significance level was reduced to 0.002 by dividing 0.05 by the number of points.
 However, these adjusted results failed to maintain statistical significance.
-Furthermore, the lack of significant differences in other channels (in this analysis) suggests that we successfully controlled the desired conditions.
+The lack of significant differences in other channels (in this analysis) suggests that the desired conditions were successfully controlled.
 These areas are typically associated with inhibition [46], emotional regulation [85], and mental workload [86], as supported by previous studies.
-Overall, our results shed light on the intricate neural dynamics of social interactions with intentional agents, offering valuable insights into the specialized roles of different prefrontal cortex channels and their adaptation to the complexities of human-AI interactions.
+Overall, the results shed light on the intricate neural dynamics of social interactions with intentional agents, offering valuable insights into the specialized roles of different prefrontal cortex channels and their adaptation to the complexities of human interactions.
 95% Confidence Interval of the Difference
@@
 -0.1090 0.2455 0.0366 -0.1827 -0.0352 -2.978 44 0.0047
-Table 2: Paired t-test Results
-Figure 8: Comparison of activation levels between human and AI interactions using the box plot of means considering standard errors.
+Table 3: Paired t-test results for oxyHb
+95% Confidence Interval of the Difference
+Mean Std. Deviation Std. Error Mean Lower Upper t df p-value
+-0.03342 0.09567 0.01426 -0.06216 -0.00467 -2.343 44 0.0237
+Table 4: Paired t-test results for deoxyHb
+Figure 9: Comparison of activation levels between human and AI interactions using the box plot of means for oxyHb considering standard errors.
+The difference in the p-value between between oxyHb and deoxyHb can be explained considering the nature of the game [87, 88, 60].
+Although both statistically indicate a difference between the two states, it seems that in this experimental setting, oxyHb demonstrates this difference more clearly.
 Discussion
-Our experiment offers valuable insights into the transition from common-sense psychology to cognitive psychology, particularly regarding interactions with intentional versus non-intentional agents.
-By utilizing fNIRS, we demonstrate the ability to capture and analyze these nuances, thereby advancing our understanding of human interactions in real-world settings and the underlying mental states involved.
-In this cognitive experiment, our primary objective is to explore the complexities of mental states elicited by interactions with intentional agents, using fNIRS to probe neural activity in ecologically valid environments.
-Our study unveils a distinctive neural signature within Channel 1, located between FpZ and AFZ according to the 10-20 system, across the frequency range of 0.00978 to 0.0196 Hz.
-Our findings are consistent with prior observations using more spatially precise neuroimaging techniques like fMRI and PET [32, 31, 36], affirming the robustness of our results across methodologies.
+The experiment offers valuable insights into the transition from common-sense psychology to cognitive psychology, particularly regarding interactions with intentional versus non-intentional agents.
+By utilizing fNIRS, the ability to capture and analyze these nuances is demonstrated, thereby advancing the understanding of human interactions in real-world settings and the underlying mental states involved.
+In this cognitive experiment, the primary objective is to explore the complexities of mental states elicited by interactions with intentional agents, using fNIRS to probe neural activity in ecologically valid environments.
+This study unveils a distinctive neural signature within Channel 1, located between FpZ and AFZ according to the 10-20 system, across the frequency range of 0.00978 to 0.0196 Hz.
+In other word, the content of thought have correlation with brain signals that captured through fNIRS imaging.
+The findings are consistent with prior observations using more spatially precise neuroimaging techniques like fMRI and PET [33, 32, 37], affirming the robustness of the results across methodologies.
 The BioImage Suite website  [^2] https://bioimagesuiteweb.github.io/webapp/ is used to generate an approximate graphical representation of the brain region illustrating in Fig.10.
@@
 This is nearly aligned with previous research that identified Brodmann area 10 and the frontal cortex as responsible for social behaviour
-Leveraging the temporal resolution advantages of fNIRS and wavelet analysis, we go deeper into the underlying brain mechanisms responsible for these observed differences.
-Our results suggest a link between this neural function and a low-frequency blood flow pattern, particularly in the mental representation related to the presence of another person.
-Notably, we found no significant difference in this fNIRS channel at higher frequencies when applying within-subject analysis, indicating that the sense of presence manifested in longer waves is associated with more stable processes.
+Leveraging the temporal resolution advantages of fNIRS and wavelet analysis, deeper exploration into the underlying brain mechanisms responsible for these observed differences is undertaken.
+The results suggest a link between this neural function and a low-frequency blood flow pattern, particularly in the mental representation related to the presence of another person.
+Notably, no significant difference in this fNIRS channel at higher frequencies was found when applying within-subject analysis, indicating that the sense of presence manifested in longer waves is associated with more stable processes.
 It appears that the representation of another person’s presence is continuous, unlike tasks involving attention or decision-making, where mental states fluctuate constantly.
-Furthermore, our straightforward approach to identifying channels related to cognitive load during the tennis game task involved imagining a binary space where the result of $Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)$ had to be positive, indicating a greater relative concentration during gameplay.
-We labeled positive results as success and negative results as failure within this binary space.
-The success-to-failure ratio for 24 data points was calculated, revealing a substantial success rate of 83% (in binomial test p < .00000001) within the frequency band of 0.0195-0.0392 Hz in Channel 1.
+Furthermore, the employed straightforward approach to identifying channels related to cognitive load during the tennis game task involved imagining a binary space where the result of $Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)$ had to be positive, indicating a greater relative concentration during gameplay.
+Positive results were labeled as success and negative results were labeled as failure within this binary space.
+The success-to-failure ratio for 24 data points was calculated, revealing a substantial success rate of 83% (in binomial test p < 0.00000001) within the frequency band of 0.0195-0.0392 Hz in Channel 1.
 This suggests that while Channel 1 is involved, different frequencies of blood flow changes are implicated in game workload and the subjective sense of interaction with intentional agents, indicating a multifunctional role for certain brain areas.
-Overall, our findings regarding the difference in frequency bands contribute to bridging the gap between common-sense psychology and scientific psychology.
-In simpler terms, it’s akin to a valve opening blood flow with a certain concentration and minor changes related to the feeling of presence, while gameplay opens another valve that transfers a concentration with more pronounced changes to the frontal cortex.
-This insight enhances our understanding of the intricate connection between the mind and the brain.
-However our results highlight a distinct difference between interactions with intentional agents and artificial agent, it’s imperative to acknowledge several limitations that may influence the interpretation of our findings as follows:
+Overall, the findings concerning differences in frequency bands serve to bridge the gap between common-sense psychology and scientific psychology.
+To simplify, consider it akin to regulating blood flow through a valve: one valve manages a consistent flow with subtle adjustments, impacting the sense of presence, while another valve oversees a flow with more pronounced changes, particularly affecting the frontal cortex during gameplay.
+Drawing on this analogy, we hypothesize that the varying speeds of different blood streams in the brain correspond to distinct mental states.
+This concept is elucidated by fNIRS, which provides both spatial and temporal resolution, further refined by computational techniques such as wavelet analysis.
+Such an approach can offer profound insights into the areas involved in various mental operations or hubs in the cerebral cortex, such as Brodmann’s Area 10 [89, 90, 91].
+Nonetheless, this remains a conjecture to be explored in future research endeavors.
+The results highlight a distinct difference between interactions with intentional agents and artificial agent.
+It’s imperative to acknowledge several limitations that may influence the interpretation of this research’s findings as follows:
   1. 1.
-    Our study’s sample population was limited to students aged between 18 and 30, potentially restricting the generalizability of our findings to other demographic groups.
+    This study’s sample population was limited to students aged between 18 and 30, potentially restricting the generalizability of the findings to other demographic groups.
   2. 2.
-    The spatial inaccuracy of fNIRS caps based on the 10-20 channel system may have impacted the precision of our results due to variations in skull shapes among individuals.
+    The spatial inaccuracy of fNIRS caps based on the 10-20 channel system may have impacted the precision of the results due to variations in skull shapes among individuals.
   3. 3.
@@
     Another important limitation of this research project was the limited number of implemented channels for measurements, while all areas of the head have the potential to receive signals.
-We focused solely on specific regions within the frontal cortex for reasons mentioned in previous sections.
+Sole focus was placed on specific regions within the frontal cortex for reasons mentioned in previous sections.
 To address these limitations and further elucidate the nuances of human-AI interactions, future research utilizing fNIRS could explore alternative methodologies and employ larger, more diverse sample populations.